OpenStudy (anonymous):

Solving Systems by Elimination. I don't know how to solve for y. I'll post my problem.

OpenStudy (anonymous):

2x + 5y = 17 6x - 5y = -9

OpenStudy (anonymous):

I know x = 1 but, how do I solve for y after plugging into x? Dividing is what I have a hard time, like when one of the numbers become a fraction...that's really confusing. :/

OpenStudy (tkhunny):

You must do better than that. Start here: 2x + 5y = 17 Solve that for y.

OpenStudy (anonymous):

Ohh. So, y = 3. I was told to do it to the 2nd equation.

OpenStudy (anonymous):

Is there a way to do it to the 2nd equation?

OpenStudy (tkhunny):

It matters not. Let's forget about x = 1 for a moment. Where did you get that, anyway? If you are going to use "Substitution"m you must pick an equatin and solve it for one f the variables. For this cause, I said, solve this for y... 2x + 5y = 17 Do that, first.

OpenStudy (anonymous):

No, not substitution. Elimination.

OpenStudy (tkhunny):

Look at that! Did you just change it or am I going blind? 2x + 5y = 17 6x - 5y = -9 It is most convenient, because we have to do nothing, to "eliminate" the y. We can just add the two equations. What do you get for that?

OpenStudy (anonymous):

Haha. It's okay, we all make mistakes. So, adding the equations would be 8x = 12?... o.o That makes no sense to me..

OpenStudy (anonymous):

Wait....no...8x = 8...... I was thinking 19 and -7. Haha!

OpenStudy (tkhunny):

That's probably because \(17 - 9 = 8 \ne 12\) Ar you happier with \(8x = 8\). That's easy to solve for x, right?

OpenStudy (anonymous):

x = 1.

OpenStudy (tkhunny):

Ding!!! That is a prize-winner. Now we have two good possibilities. Since we are trying to find a simultaneous solution, it shoudl not matter which equation we use to find y. Let's do them both, now that we know x = 1. 2x + 5y = 17 6x - 5y = -9 Substituting the known value, x = 1, we have 2(1) + 5y = 17 6(1) - 5y = -9 Solve both for y and see what you get.

OpenStudy (anonymous):

So, "y" should equal 3. Correct? :o

OpenStudy (anonymous):

Well. For the 2nd equation...no.. :/

OpenStudy (tkhunny):

Are you guessing or did you prove it?

OpenStudy (tkhunny):

6(1) - 5y = -9 6 - 5y = -9 -5y = -15 y = 3

OpenStudy (anonymous):

I proved it on scratch paper: 2(1) + 5y = 17. 2 + 5y = 17. -2 -2 ------------ 5y = 15. So y = 3.

OpenStudy (anonymous):

Ah, now I realize the 2nd equation makes more sense to me now..*Gives a Medal* :)

OpenStudy (tkhunny):

Please notice how this order of solution was arbitrary. We did x first because it looked most convenient. There is absolutely nothing wrong with finding y first. Out two equations: 2x + 5y = 17 6x - 5y = -9 Multiply the top equation by -3 -6x - 15y = -51 6x - 5y = -9 Now add the two equations. -20y = -60 And out pops y = 3. Keep your eyes open and don't get stuck in one way of thinking. Good work.

OpenStudy (anonymous):

Ahhh. Okay. Well, thank you so much! I was stressed for days on something I have not learned yet from Algebra 1. This is such a relief.